OpenStudy (johndeer34):

Suppose that y varies inversely with x, and y = 6 when x = 8. What is an equation for the inverse variation?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (johndeer34):

y=x/48?

OpenStudy (skullpatrol):

Sooo... Close!

OpenStudy (johndeer34):

y=x/48?

OpenStudy (johndeer34):

48/x i mean

OpenStudy (skullpatrol):

Yup. y = 48/x Because the form of an inverse variation is y = k/x.

OpenStudy (johndeer34):

thanks!

OpenStudy (skullpatrol):

Good job :D

OpenStudy (skullpatrol):

Begin with the definition of an inverse variation: $$\huge y=\dfrac{k}{x}$$

OpenStudy (skullpatrol):

Or $$\huge xy=k$$